Please help me solve this equation: Bill is nine years older than Jack. In Six years Bill will be twice as old as Jack. Find their present ages.
Answer by Earlsdon(6294)

Let B = Bill's present age and J = Jack's present age, then...
1) B = J+9 "Bill is nine years older than Jack."
2) B+6 = 2(J+6) "In six years Bill will be twice as old as Jack."
Substitute B = J+9 from equation 1) for the B in equation 2) to get:
(J+9)+6 = 2(J+6) Simplify.
J+15 = 2J+12 Subtract J from both sides.
15 = J+12 Subtract 12 from both sides.
3 = J
Jack's present age is 3 years and Bill's present age is 3+9 = 12 years.